OpenStudy (anonymous):

Line A is represented by the following equation: x + y = 2 What is most likely the equation for line B so the set of equations has no solution? x + 2y = 2 2x + 2y = 4 2x + y = 2 x + y = 4

OpenStudy (anonymous):

this one I really can't do haha

OpenStudy (anonymous):

D

OpenStudy (anonymous):

cool, how did you find that out? could you explain? haha

OpenStudy (daniee_bruhh):

I have no idea either lol But I trust @newhorizons

OpenStudy (anonymous):

Because then they would be parallel lines. They can also be written y = -x + 2 and y = -x + 4.

OpenStudy (anonymous):

okay thanks guys!! give me a sec to find some more questions :D

OpenStudy (anonymous):

great!

OpenStudy (anonymous):

nevermind I don't have more! thanks byee